FEL Upgrade Project
November 2001 Monthly Report

During last two weeks of the Fall User Run the FEL was used to generate approximately 100 watts in the THz region, a world record by 6 orders of magnitude; work at 5 different wavelengths obtained good data for 3 new user groups. All this was accomplished with a klystron replacement and a photocathode cesiation tossed into the operations. The machine is quite repeatable even at the 3rd harmonic of 3.15 um.

On Nov. 18th the IR Demo FEL delivered its last photons prior to being decommissioned for start of the installation process for the IR Demo Upgrade. For its last weekend of operation the IR Demo was used for four concluding runs for laser production of carbon nanotubes. The nanotube team reports that they approximately doubled their inventory of nanotube material.

Immediately after the final run, the decommissioning and dis-assembly process began. All components of the Demo were barcoded for inventory and specific components, which are going to be re-used in the Upgrade, were specifically marked. By the end of this week, all cabling in the 4F and 5F regions of the machine had been disconnecting and the backleg vacuum system had been removed. The shielding blocks were removed from the access door to the accelerator vault for removal and installation of major hardware.

As a second shift activity, the SRF team performed high power pulsed processing on the second cavity in the injector cryounit improving its maximum gradient by at least 10%. We thank Charlie Reece for organizing and overseeing this activity.

On November 8-9 the project had its third semi-annual review of the FEL upgrade project chaired by John Albertine (consultant to ONR) and John Eric (AFRL). The following is a preliminary summary of the review committee's comments. A formal report by the committee will be delivered to ONR, AFRL and DOE in December.

The committee is "very pleased and impressed with the quality of the work being done. The project addressed technical issues systematically and now all issues are well in hand. The FEL Team made a real push to get as far as they could before the end of the fiscal year (Phase 1 completion) and that is greatly appreciated. The remaining Phase 1 work (leading to the carried over cost variance) is a serious issue, but it is not large or unreasonable in high-tech project requiring R&D.

The committee doesn’t want this issue to, in any way, minimize the excellent work and progress made. Jefferson Lab’s success and progress is largely responsible for the credibility FELs are beginning to enjoy.

The Navy desired the upgrade to operate at 10 kW at 1-2 micron range, and there will be a request from ONR to DOE coming to ask what is the easiest way to get there, the team should be thinking of what it would take."

Following the November semiannual project review, we are taking the committee's advice and carefully analyzing all budget categories for the remaining tasks in the Upgrade project to identify options for cost savings or scope reduction. These steps are also necessary since it appears highly likely that the FY02 DOD budget allocations (that could provide some relief to our tight budget situation) will be delayed well into the fiscal year.

The subcontract for participation by the Aerospace Corporation in the UV FEL project was forwarded to the DOE Site Office for approval.

On Nov. 29-30, we had a visiting team from AES, Inc. for discussions of the joint work statement for the start-up of the 100 mA injector project. This start-up is possible because AES was recently awarded a Phase 2 SBIR from SMDC (with supporting funds from NAVSEA) to design and fabricate key components of a 750 MHz injector cryounit compatible with the JLab FEL photogun. We also had visitors in November from the Indian Head NSWC (William Walker and Walter Ott) who are interested in potential collaboration with JLab and SURA universities, from LBNL to discuss potential collaboration on next generation x-ray sources; and visitors from Dahlgren Labs and Northrop Grumman to discuss cryoplant designs for high average power FELs.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k. The first phase of the project was fully obligated as of the end of September. The total Phase 1 project cost to date-actual costs plus commitments is $9,029k as planned for Phase 1 completion. Costs accrued to Phase 1 during this period are bookings for previously committed funds as equipment is received. When the open commitments are fully costed on or before December 30, 2001, we predict that the remaining Phase 1 cost and schedule variances will be approximately equal to -$1,300k. Accordingly, the final Phase 1 cost performance report is planned for January 2002. We expect to recover remaining Phase 1 cost and schedule variances by the completion of the Phase 2 barring unforeseen difficulties.

The Phase 2 project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002. Work for Phase 2 proceeded to plan during this reporting period. Actual charges of $574k were accrued for November for total accrued costs of $1,508k to date for Phase 2. The bulk of this effort involves long term procurements of components designed during Phase 1. No significant earned value is planned for the first quarter of FY02. We expect to meet the planned cost at completion of Phase 2 effort. At present the cost variance for Phase 2 is -$1,241k and the schedule variance is $13k. The program is presently on track to begin commissioning efforts (presently un-funded) in September 2002.

The UV FEL project was approved for $2,735k for a performance period from September 2001 to September 30, 2002. Efforts were initiated in October with the official project start toward the project deliverables: design of the optical cavity system, design of the UV FEL electron beam and optical transport systems, and production of harmonically generated UV light. Significant earned value credit should be available by the end of the first quarter of FY02 for all of these activities.

WBS 4 (Injector):
By the end of the Fall Run on Nov. 19, the current GaAs wafer in the gun has delivered over 5300 Coulombs. We plan to give it a comfortable retirement.

Modifications made to the cathode protection door assembly were checked and seemed to work much better, some parts are out for silver plating. The electrical wireways for the cesium channels and charge collection ring were submitted to the shop. The three getter pump assemblies for the gun chamber were received. The parts for the gun chamber have been completed by the shop; they will be cleaned prior to welding.

We baked the Field Emission Test (FET) system with an implanted electrode installed; testing will start in December.

WBS 5 (SRF):
The 1st upgrade thermal shield has been received from the vendor. Inspection noted a few minor discrepancies that will be fixed in house. The 1st and 2nd upgrade end cans have been received. Preliminary inspection completed with final inspection in January. A decision was made to redesign the helium vessel. Conceptual work completed and detailed design is in process.

WBS 6 (RF):
Zone 2 - The 5 watt RF Amplifiers were received, but are not yet installed and tested.

Zone 3 - Essentially complete. Missing Kirk Key lock on HVPS and testing of arc & IR detectors. These will be completed after Fall User run.

Zone 4 - Essentially complete. Missing power supply for 5 watt RF amplifiers and system test of position 8 klystron & circulator. Will be completed after the Fall User Run run.

Master Oscillator - Equipment for zones 3 & 4 is in-house, but can not be installed until after the Fall OPS run.

100 kW Klystron - The LCW plumbing of the klystron cart is nearly finished. The flow rates will be checked with the existing LCW taps to determine the power levels for preliminary testing of the klystrons.

New ion pump controllers for the klystrons were received and connected.

Quarter HVPS - The circuit design of the HVPS is nearly complete. Capacitors were ordered for the DC

filter section. Resistors will be specified shortly. The schematic will be updated and an informal review held.

Quarter Cryounit - High Peak Power Processing (HPPP) was run on cavity 3. The field emission of the cavity was improved. Additional processing will be performed today and several days next week at higher power levels.

WBS 8 (Instrumentation):
Design for the VME Interface board is complete and the Digital Isolation Module is in the "layout" phase. We have received additional Four Channel Timing/Logic Boards and will begin installing parts and testing after the holiday.

A barcode tracking system for IR Demo FEL removal activities has been developed. All major components in the current machine will be barcoded for tracking purposes prior to their removal from the vault. A database for all components is being designed to be implement as new installations take place.

Within one week after the termination of the Fall User Run on Nov. 19, all cabling was removed from the 4F and 5F regions in support for removing the components in these areas.

Motion sensor cameras and time-lapse VCRs will be installed to track the upgrade process on video.

The individual user labs are being re-organized and re-stocked for storage purposes until they will be used again.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
• We inspected the initial winding of the first coil as well as bend up fixtures and the potting fixture machining at Magnet
Enterprises International. We have enlisted the help of an epoxy and coil failure expert, Jeffery Hubrig of Innovation Services to observe and comment on the coil potting.
Injector Dipoles (DU/DV)
• DULY Research is almost finished revising small portions of the final assemblies. The specifications are being written to match the drawings and revised with what we have learned from our procurements of magnets to date.
Arc Dipoles (GY, GX, GQ)
• The detail drawings of the bend and Reverse Bend Dipoles (GX & GQ) are finished except for a adjustment to the height of the pole tip to increase the width of the good field region.
• The magnetic model of the above dipoles is being exercised. The GX is nearly through and the GQ is starting. The drawings will be revised when the adjustment is over.
• The bids for the 180° Dipoles (DY) are due on the first working day of December.
Magnet Measurement
• The drawings for the measurement stands for the two quadrupoles are being made.
• The parts for the dipole measurement stand are signed off and in procurement.
• We continued work on the probe that will measure the sextupole and the octupole.
Multipoles and Correctors
3 inch Quad (QX)
• The sample aluminum version of the core pole was machined and passed inspection before finish machining (except for the 45° chamfer).
• The core vender, New England Techni-Coil is making the final adjustment to the 45 degree chamfer on the pole edge of the first article core.
• New England Techni-Coil, the assembly vendor fed back to us that the coils were not potted in the style he was anticipating. After some exchanging of photographs and discussions with their manufacturer, Magnet Enterprises International, the question was resolved. We sent Kevin Sullivan to inspect the coils and consult with New England Techni Coil about the details of assembly.
• Inspection there seemed to indicate turn to turn shorts in the coils. The controversy was resolved by unwrapping the copper tubing of the cooling circuits. The wrapped configuration was giving false indications of shorts.
• A second potential problem is being resolved by a thermal test. Sporadic brazing of the tubes to cooling plates on the initial series of coils is being compared in a thermal test with coils that have continuous brazing.
• A shipment of 64 fully qualified coils is being readied at Magnet Enterprises International.
Trim Quad (QT)
• After the discussions with venders about our QX coils during the procurement process, Tom Hiatt determined items he wanted to change in the manufacturing specification. After the revision, we went out with the final bid cycle. The bids were received and evaluated. The low bidder is being asked to detail a cost breakdown because of the spread in the quotations.
Sextupole (SF)
• We composed a statement of work for DULY Research to model the sextupole. They responded and are now performing the modeling.
Octupole (OT)
• Work on this magnet is on furlough until completion of the sextupole.
Corrector Dipoles (DB, DJ)
• Milhous Co. is about 6 shy of delivering the entire order DB and DJ coils.
Beam Line and Vacuum
• Dave Waldman completed the task order for design of the Arc Chambers utilizing the Engineering Services contract with Advanced Energy Systems. He chaired the kick-off meeting to start the task order with AES.
• Several sample chambers with the thicker plates necessary for the upgrade have been made to test out the magnetic properties.
• Design work on girders and stands continued. All details of the region between the cryomodules are now complete.
• All return leg girders have been received.
• All the new three leg pedestals are now out for bid.

WBS 10 (Wiggler):
We cleaned up some punch list items from the vacuum chamber fit up exercise and began reinstallation of the measurement apparatus to take a few runs to clear up some minor questions concerning the measurement sets taken to date.

WBS 11 (Optics):
With assistance from K. Capek (Accel. Div.), an ultraviewer was assembled and is undergoing tests in User Lab 6. We have "mocked up" one-half of the optical cavity beam path in User Lab 2 and confirmed calculations of beam sizes at various points in the path. This is an important check, as we are close to releasing drawings for the OCMMS (optical cavity mirror metrology system) beamline components. Before the Thanksgiving break the OCMMS drawings were reviewed, marked up, and returned for revision

Detailing continues on the mounting of the optical cavity components to the vacuum vessel.

We have received the following parts for the optical cavities: ball bearings for the leadscrew, flex pivots for the yaw axes, and flex pivots for the pitch axes. The bearings are being cleaned for UHV service.

Detailing concentrated on components from the "big C" (the yaw axis yoke) through the mirror holders. We have identified a vendor for sputter-coating the leadscrews and are going out for procurement.

Meetings continued with sales representatives of ultrafast lasers. This laser will be used as a FEL cavity dynamical diagnostic, and for diagnostics of the e-beam, using EO sampling. Further meetings were held on procurement of a new drive laser.

The first week of November (5-9) was devoted to lasing near 1 micron. Operations were delayed somewhat by the loss of an injector klystron on Monday and the need to cesiate the cathode on Thursday. Despite this, Jacob Grun from NRL got all the data he needed on LOP-01-049, a thermal annealing of silicon. Harvey Rutt, working on LOP-01-044, managed to create what, to all appearances, look like high quality zinc chalcogenide films using ultrafast laser ablation with the frequency doubled FEL radiation at 525 nm. In parallel with Harvey Rutt's work we worked on LOP-01-057--ablation rates at short wavelengths. Once the injector RF was repaired, the machine has been quite reliable this week, though the quantum efficiency was initially low.

The last week of the Fall User Run (Nov. 12-16) was devoted to THz studies and operation near 1.6 microns using a new mirror set. The laser lased well at 1.6 microns and beam was provided for LOP-01-043--Absorption measurements in GaLaS glasses--performed by Harry Rutt. Jim Boyce ran parasitically to try to see doubly diffracted X-rays. This turned out to be very difficult due to frequent changes in the system. The detection of doubly diffracted X-rays requires many hours of steady electron beam and lasing. In the end, no clear signal was seen. Visitors from Brookhaven National Lab and Lawrence Berkeley National Lab gathered excellent data on the spectrum and power of the THz radiation from two different ports on the machine. To demonstrate how strong this radiation is, we lit a match using the THz radiation. Efforts by Masahiko Tani from Japan to image using the THz were not successful due to alignment difficulties. Many thanks go out to the many people who worked heroically on this experiment to get a signal for Prof. Tani.

On Friday, Nov. 16, mirrors were switched from 1.6 micron operation to 3 micron operation. This allowed the nanotube collaboration to try some new targets on a special last run of the IR Demo on Nov. 17-18.